OpenStudy (anonymous):

Use the distributive property to simplify. 3q + 2(q + 1)

OpenStudy (anonymous):

@XxPikachuxX 3q + 2(q + 1) = 3q + (2)(q) + (2)(1) = 3q + 2q + 2 = 3q + 2q + 2 = (3q + 2q + 2 = (3q + 2q) + (2) = 5q + 2

OpenStudy (anonymous):

3q + 2*(q + 1) 3q + 2*q + 2*1 3q + 2q + 2 5q + 1

OpenStudy (anonymous):

oops 5q + 2, typo